Here few guidelines on how to do clean chroma decay measurements in case you have the occasion:
- As soon as you arrive at injection play the task "STOP FIDEL TRIMMING" (contained in LHC PILOT SEQUENCE -> PREPARE LHC FOR INJECTION)
- Revert the B3 correction trims to the last "FIDEL trim at time 0.0 sec". This will cancel any existing (or trimmed before switching FIDEL off) correction.
- Set the chroma knob to:
- -1 H
- -10 V
- This should give you a chroma:
- ~0 H
- ~20/25 V
- Switch the RF modulation ON (DELTA_f = 100 Hz)
- If the chroma differs from the values above please adjust it
- Leave it for about an hour
The measurements should be done as close as possible to the beginning of the injection plateau.
For FIDEL parametrization it would be very good to have (at least) 2 sets of measurements with different powering history, namely after:
- Precycle (T_FT = 600 sec)
- Ramp down after long(er) flattop